EOQ for Multiple Products: In another location, the distributor of the previous problem stocks four different...

  1. EOQ for Multiple Products: In another location, the distributor of the previous problem stocks four different items in common warehouse space. Each item is described by an annual demand rate, a fixed cost per order, a holding cost per year, a unit purchase cost, and a space requirement. The data in the following table describe the four products.

Item

1

2

3

4

Demand

5000

10,000

30,000

300

Fixed cost

400

700

100

250

Holding cost

50

25

8

100

Purchase cost

500

250

80

1000

Space (sq. ft)

12

25

5

10

  1. Create an Excel Spreadsheet with Solver to answer the following questions.
  2. Considering each product separately, as if it were independent from the others, what are the respective economic order quantities? What is the total annual cost of ordering, holding, and purchasing across the four products at these order quantities?
  3. What is the minimum total annual cost for the four products if the average space taken up must be no more than 12,000 sq. ft?
  4. What is the minimum total annual cost for the four products if the average space taken up must be no more than 12,000 sq. ft and the number of orders per year must be no more than 65?
  5. In part (b), with a square-foot limit on storage space, what is the economic value of more space?
  6. In part (c), what is the economic value of more space?
Answer #1

a) Excel Spreadsheet model is following:

Item 1 2 3 4
Demand (D) 5000 10000 30000 300
Fixed cost (A) 400 700 100 250
Holding cost (H) 50 25 8 100
Purchase cost (C) 500 250 80 1000
Space (sq. ft) (s) 12 25 5 10
Order quantity (Q) =SQRT(2*B2*B3/B4) =SQRT(2*C2*C3/C4) =SQRT(2*D2*D3/D4) =SQRT(2*E2*E3/E4)
Total space
Average space (s*Q/2) =B6*B8/2 =C6*C8/2 =D6*D8/2 =E6*E8/2 =SUM(B10:E10)
# of orders
Number of orders =D/Q =B2/B8 =C2/C8 =D2/D8 =E2/E8 =SUM(B12:E12)
Ordering cost =A*D/Q =B3*B12 =C3*C12 =D3*D12 =E3*E12
Holding cost = H*Q/2 =B4*B8/2 =C4*C8/2 =D4*D8/2 =E4*E8/2
Purchasing cost =D*C =B2*B5 =C2*C5 =D2*D5 =E2*E5
Total Cost
Total cost = =SUM(B12:B15) =SUM(C12:C15) =SUM(D12:D15) =SUM(E12:E15) =SUM(B17:E17)


b) Calculation of EOQ, annual cost of ordering, holding and purchasing costs is shown below.

c) Minimum cost ordering policy is determined by Solver as follows:

Order quantity of item 1 = 270

Order quantity of item 2 = 657

Order quantity of item 3 = 798

Order quantity of item 4 = 35

d) The optimal solution is following:

e)

With limitation on storage space, increase in total cost = 7743954-7743725 = $ 229

Without limitation, total additional space used = 13410-12000 = 1410 sqft

So, economic value of more space = 229/1410 = $ 0.16 per sqft